OpenStudy (anonymous):

solve the fallowing systems of equations algebraically Y=7x-8 8x-2y=1

OpenStudy (anonymous):

8x-2(7x-8)=1 8x-14x+16=1 -6x=-15 x=15/6 Y= 7(15/6)-8 Y=19/2 8(15/6)-2(19/2)=1 20-19=1